# Rotationally symmetric translating solitons of fully nonlinear extrinsic geometric flows: Classification and Applications

### v1

@@ 1. Introduction
Translating solitons of extrinsic geometric flows whose normal velocity is given by a positive curvature function arise as canonical models for type II singularities, where curvature concentration behaves more irregularly than for spheres and cylinders during their deformation, see [chow2020extrinsic, lynch2023differential]. Moreover, they also serve as building blocks for the long-term behavior of these extrinsic curvature flows, see [rengaswami2023ancient].
@@
The first goal of this article is to develop a unified framework for rotationally symmetric translators of $\gamma$-flows, where $\gamma(\lambda)$ is a curvature function which it is also the normal velocity of a family of evolving hypersurfaces. The class of curvature functions that we are interested in this article are positive or signed, $\alpha$-homogeneous and symmetric function that are increasing in each variable when they are evaluated in the principal curvatures of this family of hypersurfaces. Then, by using an implicit formulation of the translator equation in terms of the functions $x=g_{\pm}(y,\pm 1)$ associated with the level sets of $\tilde{\gamma}(x,y,\ldots,y)=\pm 1$, we establish the fine asymptotic expansion at infinity of bowl-type translators, see Theorems 4.1 and 4.8.
Our second main contribution is the construction and classification of rotationally symmetric catenoidal translators for suitable signed curvature functions (Section 5). For each $R>0$, we construct the catenoidal translator $W_{R}$ as a hypersurface of revolution generated by a curve with a single neck of radius $R$, and we show that, after removing a large ball, $W_{R}$ possesses two branches given by the graph of a function. Furthermore, we prove that its upper branch has the same bowl-type asymptotics as in the positive curvature case, while the lower branch exhibits either a bowl-type end or a different growth regime, depending on the behavior of $g_{-}(\cdot,-1)$ near the origin, see Theorem 5.1 and the pictures below.
@@ 2. Preliminaries
Firstly, given an immersed hypersurface $\Sigma_{0}=F_{0}(\Sigma)\subset{\mathbb{R}}^{n+1}$, which we assume smooth throughout the article unless stated otherwise, we recall some basic notions of extrinsic geometry in ${\mathbb{R}}^{n+1}$. The principal curvatures $\lambda(p)=(\lambda_{1}(p),\ldots,\lambda_{n}(p))$ of $p\in\Sigma_{0}$ are the eigenvalues of the shape operator $\mathcal{W}:T_{p}\Sigma_{0}\to T_{p}\Sigma_{0}$, defined at $p\in\Sigma_{0}$ by $\mathcal{W}(\vec{X})=-\nabla_{\vec{X}}\vec{\nu}$, where $\nu$ is a unit normal vector with a fixed orientation in ${\mathbb{R}}^{n+1}$ and $\nabla$ is the Euclidean connection.
@@ 2.1. The class of $\gamma$-flows
Next we specify the class of curvature functions $\gamma$ that we will work with.
@@
We emphasize that the hypotheses defining the class of positive, 1-homogeneous curvature functions, together with second-order structural properties such as convexity, concavity, and inverse concavity, supported on the positive cone $\Gamma_{+}$, guarantee short-time existence and uniqueness [^2] Up to tangential diffeomorphisms of $\Sigma_{t}$. for strictly convex ($\lambda\in\Gamma_{+}$) compact initial data without boundary, see [chow2020extrinsic].
@@
    There is an open cone $\tilde{\Gamma}:=\{(x,y)\in\mathbb{R}^{2}:(x,y,\ldots,y)\in\Gamma\}$ and a continuous extension $\tilde{\gamma}:\tilde{\Gamma}_{+}\cup\tilde{\Gamma}_{-}\to\mathbb{R}$, where
@@ 2.2. Translators
This definition is motivated by the behavior of the $\gamma$-flow on round generalized cylinder of the form $\mathbb{S}^{n-p}\times\mathbb{R}^{p}$ in ${\mathbb{R}}^{n+1}$. Under $k$-nondegeneracy, cylinders with $p<k$ contract in finite time to a $k$-hyperplane, whereas under $k$-degeneracy, cylinders with $p\geq k$ are stationary solutions of the $\gamma$-flow.
We now list examples of translating solitons for general curvature functions:
  • $\alpha$-grim-reaper cylinders. For (n1)(n-1)-nondegenerate, normalized $\alpha$-homogeneous positive curvature functions (i.e., $\gamma(0,\ldots,0,1)=1$), the family of translating cylinders, unique up to scaling and rotation, isometric to $C\times\mathbb{R}^{n-1}$, where $C$ is a one-dimensional $\alpha$-grim-reaper solving $\lambda^{\alpha}=\langle\nu,e_{2}\rangle$, see [bourni2022ancient, chou2001curve]. These cylinders are geodesically complete for $\alpha\leq 1$. Moreover, for $\alpha\in\left(\tfrac{1}{2},1\right]$, the generating curves are bounded on a finite interval, so the cylinders lie in slabs of finite width; whereas for $\alpha\in(0,\tfrac{1}{2})$, they are defined in all $\mathbb{R}$, yielding entire graphical solutions.
@@
  • Catenoidal translators. This family consists of non-convex rotationally symmetric hypersurfaces parameterized by their positive distance to the translation axis, and they are singular along the plane $\{x_{n+1}=0\}$ in a large class of ambient spaces. These solutions are currently known only for the elementary symmetric polynomials, i.e., $\gamma=S_{k}$. As in the bowl-type case, we encounter a dichotomy: for $k=1,\ldots,n-1$, the catenoidal translators are rotational bi-graphs over the complement of a ball where each branch is asymptotic to a vertical translate of the bowl-type soliton [clutterbuck2007stability]. For $k=n$, the catenoidal family instead exhibits a conical singularity; we refer to [lima2025translators] for further details on the classification of these solutions .
We also remark that in the mean curvature case $\gamma=H$, there is a large variety of translating solitons, including the $\Delta$-wings family and several semi-graphical translators constructed and classified in [martin2024classification].
@@ 2.3. Rotationally symmetric translators
We consider $G:\tilde{\Gamma}\times(0,\infty)\to{\mathbb{R}}$ defined by $G(x,y,z)=\tilde{\gamma}(x,y)-z$. We note the following:
@@
As an example, we can compute $g_{\pm}$ explicitly for the $1$-homogeneous normalized Hessian quotient curvature functions
@@
3. Barrier methods
In this section we develop the barrier method for rotationally symmetric translators whose generating curve $\mathtt{c}(s)$ is given as a vertical graph $(r,u(r))$ in $\mathbb{R}^{2}$.
@@ 3. Barrier methods
The barrier method for (5) combines a comparison principle with suitable barriers that ensure the implicit functions $g_{\pm}(y,\pm 1)$ are well defined. It also provides first-order control for analyzing the behavior of solutions.
@@
We note that $g_{+}(y,1)$ extends continuously to the endpoints of $U_{+}$ by using (4). Indeed,
@@
As an example of a curvature function that satisfies the hypotheses of this proposition, we consider $\gamma_{k}=Q_{k}=\frac{S_{k}}{S_{k-1}}$. In this case $g_{-}(y,-1)=\frac{\binom{n-1}{k}y\left(-1-y\right)}{\left(\binom{n-1}{k-1}y+\frac{\binom{n-1}{k}\binom{n-1}{k-2}}{\binom{n-1}{k-1}}\right)}$, and we notice that $\partial_{y}g_{-}(0,-1)=\frac{-(n-k+1)}{k-1}$. Since this number indicates the exponent in the growth order, we will observe in the following sections that for this curvature function the lower branch of the catenoidal translator $W_{R}$, has a growth order of $\frac{-a(R)(k-1)}{2(k+1)-n}r^{\frac{2(k+1)-n}{k-1}}$, which is negative when $k<\frac{n-2}{2}$, zero when $k=\frac{n-2}{2}$ is an integer, and positive otherwise.
@@ 4. Fine Asymptotic of Bowl-type Solutions
The first step was established in [rengaswami2024classification] as part of our bootstrapping scheme.
@@ A Taylor expansion of $g$ at $y=1$ gives
for $1$-degenerate $\alpha$-homogeneous positive curvature functions. For instance, a condition ensuring the existence of entire bowl-type solutions for $1$-degenerate curvature functions is $\alpha\in\left(\frac{1}{3},\frac{1}{2}\right)$, or $g_{+}(y,1)\to 0$ with decaying growth $g_{+}(y,1)=O\!\left(y^{1-2\alpha}\right)$ as $y\to\infty$.
@@
Let $\gamma:\overline{\Gamma}\to[0,\infty)$ be a $1$-degenerate $\alpha$-homogeneous positive curvature function with $k_{\gamma}>3\alpha-1$, and $\alpha\geq 1$. Then any solution to (15) behaves as $v=w+o(w^{-1})$ as $r\to\infty$ where $w(r)=A_{\gamma}r^{d_{\gamma}}$ and $d_{\gamma}=\frac{\alpha(k_{\gamma}+1)}{k_{\gamma}-2\alpha+1}$, $A_{\gamma}=\left(\frac{d_{\gamma}}{c_{\gamma}}\right)^{\frac{\alpha}{2\alpha-1-k_{\gamma}}}$, $k_{\gamma}=\min\{k\geq 1:\ c_{k}\neq 0\}$, where $c_{k}$ is the coefficient in the Laurent expansion at infinity of the implicit function $x=g_{+}(y,1)$, and $c_{\gamma}=c_{k_{\gamma}}>0$.
@@
As in Section 4, we divide the proof of Theorem 4.8 into three propositions.
@@
The preceding proposition completes the proof of Theorem 4.8.

### v2

@@ 1. Introduction
Translating solitons of extrinsic geometric flows whose normal velocity is given by a positive curvature function arise as canonical models for eternal evolutions and, for certain classes of curvature functions, type II singularities, where curvature concentration behaves more irregularly than for spheres and cylinders during their deformation along the singularity formation, see [chow2020extrinsic, lynch2023differential]. Moreover, they also serve as building blocks for the long-term evolution, such as ancient evolutions, of these extrinsic curvature flows, see [rengaswami2023ancient].
@@
The first goal of this article is to develop a unified framework for rotationally symmetric translators of $\gamma$-flows, where $\gamma(\lambda)$ is a curvature function which it is also the normal velocity of a family of evolving hypersurfaces.
The class of curvature functions that we are interested in this article consists of positive or signed, $\alpha$-homogeneous, symmetric functions that are increasing in each variable supported in open cones containing the positive cone $\Gamma_{+}=\left\{\lambda\in\mathbb{R}^{n}:\lambda_{i}>0\right\}$, when they are evaluated at the principal curvatures of this family of hypersurfaces. Then, by using an implicit formulation of the translator equation in terms of the functions $x=g_{\pm}(y,\pm 1)$ associated with the level sets of $\tilde{\gamma}(x,y,\ldots,y)=\pm 1$, we extend the fine asymptotic expansion at infinity of bowl-type translators given in [cogo2023rotational] to general positive $\alpha>\frac{1}{3}$ homogeneities and $1$-degeneracy order; see Theorem 4.1 for the $1$-nondegenerat and Theorem 4.8 for the $1$-degenerate.
Our second main contribution is the construction and classification of rotationally symmetric catenoidal translators for suitable signed curvature functions (Section 5). For each $R>0$, we construct the catenoidal translator $W_{R}$ as a hypersurface of revolution generated by a curve with a single neck of radius $R$, and we show that, after removing a large ball, $W_{R}$ possesses two branches given by the graph of a function. Furthermore, we prove that its upper branch has the same bowl-type asymptotics as in the positive curvature case, while the lower branch exhibits either a bowl-type end or a different growth regime, depending on the first order behavior of $g_{-}(\cdot,-1)$ near the origin, see Theorem 5.1 and the pictures below.
@@ 2. Preliminaries
Firstly, given an immersed hypersurface $\Sigma_{0}=F_{0}(\Sigma)\subset{\mathbb{R}}^{n+1}$, which we assume smooth throughout the article unless stated otherwise, we recall some basic notions of extrinsic geometry in ${\mathbb{R}}^{n+1}$. The principal curvatures $\lambda(p)=(\lambda_{1}(p),\ldots,\lambda_{n}(p))$ of $p\in\Sigma_{0}$ are the eigenvalues of the shape operator $\mathcal{W}:T_{p}\Sigma_{0}\to T_{p}\Sigma_{0}$, defined at $p\in\Sigma_{0}$ by $\mathcal{W}(\vec{X})=-(\nabla_{\vec{X}}\vec{\nu})^{\top}$, where $\nu$ is a unit normal vector with a fixed orientation in ${\mathbb{R}}^{n+1}$, $\nabla$ is the Euclidean connection and $(\cdot)^{\top}$ is the orthogonal projectio to $T_{p}\Sigma_{0}$.
@@ 2.1. The class of $\gamma$-flows
Next we specify the class of curvature functions that we will work with.
@@
We emphasize that the hypotheses defining the class of positive, 1-homogeneous curvature functions, together with second-order structural properties such as convexity, concavity, and inverse concavity, supported on the positive cones $\Gamma_{+}^{n}$, guarantee short-time existence and uniqueness [^2] Up to tangential diffeomorphisms of $\Sigma_{t}$. for strictly convex ($\lambda\in\Gamma_{+}$) compact initial data without boundary, see [chow2020extrinsic].
@@
    There is an open cone $\tilde{\Gamma}=\{(x,y)\in\mathbb{R}^{2}:(x,y,\ldots,y)\in\Gamma\}$ and a continuous extension $\tilde{\gamma}:\tilde{\Gamma}_{+}\cup\tilde{\Gamma}_{-}\to\mathbb{R}$, where
@@ 2.2. Translators
This definition is motivated by the behavior of the $\gamma$-flow on round generalized cylinder of the form $\mathbb{S}^{n-p}\times\mathbb{R}^{p}$ in ${\mathbb{R}}^{n+1}$. Under $k$-nondegeneracy, generalized cylinders with $p\leq k$ contract in finite time to a $k$-hyperplane, whereas under $k$-degeneracy, cylinders are stationary solutions of the $\gamma$-flow.
We now list examples of translating solitons for this class of curvature functions:
  • Nondegenrate $\alpha$-grim-reaper cylinders. For (n1)(n-1)-nondegenerate, normalized $\alpha$-homogeneous positive curvature functions (i.e., $\gamma(0,\ldots,0,1)=1$), the family of translating cylinders, unique up to scaling and rotation, isometric to $C\times\mathbb{R}^{n-1}$, where $C$ is a one-dimensional $\alpha$-grim-reaper solving $\lambda^{\alpha}=\langle\nu,e_{2}\rangle$, see [bourni2022ancient, chou2001curve]. These cylinders are geodesically complete for $\alpha\leq 1$. Moreover, for $\alpha\in\left(\tfrac{1}{2},1\right]$, the generating curves are bounded on a finite interval, so the cylinders lie in slabs of finite width; whereas for $\alpha\in(0,\tfrac{1}{2})$, they are defined in all $\mathbb{R}$, yielding entire graphical solutions.
  • Degenrate $\alpha$-grim-reaper cylinders. When $\gamma=S_{n}=K$ is the (non-normalized) Gaussian curvature, the associated flow is the $n$-MCF, and $K$ is an (n1)(n-1)-degenerate positive curvature function in our sense. In this setting, there exist grim-reaper-type translators in both $\mathbb{R}^{n+1}$, constructed as vertical graphs whose level sets are parallel umbilical hypersurfaces and whose height functions solve a first-order ODE depending only on the distance parameter; see [lima2025translators]. These examples are complete, have unbounded height, and are asymptotic to vertical hyperplanes, providing canonical one-dimensional model profiles for the $K$-flow analogous to the classical grim reaper for $H$-flow.
@@
  • Catenoidal translators. This family consists of non-convex rotationally symmetric hypersurfaces, parameterized by their positive distance to the translation axis, and they are singular along the plane $\{x_{n+1}=0\}$ in a large class of ambient spaces. These solutions are currently known only for the elementary symmetric polynomials, i.e., $\gamma=S_{k}$. As in the bowl-type case, we encounter a dichotomy: for $k=1,\ldots,n-1$, the catenoidal translators are rotational bi-graphs over the complement of a ball, where each branch is asymptotic to a vertical translate of the bowl-type soliton [clutterbuck2007stability, lima2025translators]. For $\gamma=S_{n}=K$, de Lima and Pipoli [lima2025translators] construct catenoid-type rotational translators to $n$-MCF which are non-convex and arise as vertical graphs over the complement of a ball (or over annular regions), with level sets given by parallel umbilical hypersurfaces. These $K$-catenoids are properly embedded, typically have two ends, and each end is asymptotic to a vertical translate of a $K$-bowl-type translator, thus mirroring the behavior of the classical translating catenoids in the mean curvature case. In some dimensions, they exhibit mild $C^{2}$-singularities along a central sphere or cone, while still qualifying as genuine translators in the sense that the translator equation $K=\langle\nu,e_{n+1}\rangle$ holds away from a set of null measure.
We also remark that in the mean curvature case $\gamma=H$, there is a large variety of translating solitons, including the $\Delta$-wings and the annuloid family [hoffman2019graphical, bourni2020existence, HMWS-deltawings]. In addition, there exist numerous complete convex non-rotationally symmetric examples [wang2011convex]. More recently, semigraphical translators have been constructed and classified [HMWS-tridents], with corresponding uniqueness results [martin2024classification]. For a detailed overview of these developments, we refer the reader to the survey [HIMW-notes].
@@ 2.3. Rotationally symmetric translators
We consider $G:\tilde{\Gamma}\times(0,\infty)\to{\mathbb{R}}$ defined by $G(x,y,z)=\tilde{\gamma}(x,y)-z$, to observe the following:
@@
Remark 2.8.
This was the barrier method (see next section) used to rigorously construct bowl-type solutions in [rengaswami2024classification] since its has a removable singularity in their setting at $r=0$.
@@
As an example, we compute $g_{\pm}(y,z)$ explicitly for the $1$-homogeneous normalized Hessian quotient curvature functions
@@
3. Barriers method
In this section we develop the barriers method for rotationally symmetric translators whose generating curve $\mathtt{c}(s)$ is given as a vertical graph $(r,u(r))$ in $\mathbb{R}^{2}$.
@@ 3. Barriers method
The barriers method for (5) combines a comparison principle with suitable barriers that ensure the implicit functions $g_{\pm}(y,\pm 1)$ are well defined. It also provides first-order control for analyzing the behavior of their solutions.
@@
We note that $g_{+}(y,1)$ extends continuously to the endpoints of $U_{+}$ by using (4). Indeed, we have the following equations
@@
As an example of a curvature function that satisfies the hypotheses of the above proposition, we consider $\gamma_{k}=Q_{k}=\frac{S_{k}}{S_{k-1}}$. In this case $g_{-}(y,-1)=\frac{\binom{n-1}{k}y\left(-1-y\right)}{\left(\binom{n-1}{k-1}y+\frac{\binom{n-1}{k}\binom{n-1}{k-2}}{\binom{n-1}{k-1}}\right)}$, and we notice that $\partial_{y}g_{-}(0,-1)=\frac{-(n-k+1)}{k-1}$. Then, the lower branch of the catenoidal translator $W_{R}$ from the Section 5, has a growth order of $\frac{-a(R)(k-1)}{2(k+1)-n}r^{\frac{2(k+1)-n}{k-1}}$, which is negative when $k<\frac{n-2}{2}$, zero when $k=\frac{n-2}{2}$ is an integer, and positive otherwise.
@@ 4. Fine Asymptotic of Bowl-type Solutions
The first step was established in [rengaswami2024classification]; we now extend this to second-order expansion and more general homogeneous curvature speeds.
@@ A Taylor expansion of $g$ at $y=1$ gives
for $1$-degenerate $\alpha$-homogeneous positive curvature functions. For instance, a condition ensuring the existence of entire bowl-type solutions for $1$-degenerate curvature functions are $\alpha>\frac{1}{2}$ and $g_{+}(y,1)\to 0$ with decaying growth $g_{+}(y,1)=O\!\left(y^{1-2\alpha}\right)$ as $y\to\infty$.
@@
Let $\gamma:\overline{\Gamma}\to[0,\infty)$ be a $1$-degenerate $\alpha$-homogeneous positive curvature function with $\alpha\geq 1$. Then any solution to (15) behaves as
  • If $k_{\gamma}>3\alpha-1$, then $v=w+o(w^{-1})$ as $r\to\infty$ where $w=A_{\gamma}r^{d_{\gamma}}$ and $d_{\gamma}=\frac{\alpha(k_{\gamma}+1)}{k_{\gamma}-2\alpha+1}$, $A_{\gamma}=\left(\frac{d_{\gamma}}{c_{\gamma}}\right)^{\frac{\alpha}{2\alpha-1-k_{\gamma}}}$, $k_{\gamma}=\min\{k\geq 1:\ c_{k}\neq 0\}$, where $c_{k}$ is the coefficient in the Laurent expansion at infinity of the implicit function $x=g_{+}(y,1)$, and $c_{\gamma}=c_{k_{\gamma}}>0$.
  • If $k_{\gamma}=2\alpha-1$, then $v=e^{c_{\gamma}\frac{r^{2\alpha}}{2\alpha}+o(r^{2\alpha})}$, as $r\to\infty$
@@
As in Section 4, we divide the proof of Theorem 4.8 into several propositions.
@@
The preceding proposition completes the proof of Theorem 4.8 in the case $k_{\gamma}>2\alpha-1$. Now, given $\varepsilon>0$, we continue with the case $k_{\gamma}=2\alpha-1$ and consider the solution $w_{\varepsilon}(r)$ of
$\begin{cases}w_{\varepsilon,\pm}^{\prime}=(1\pm\varepsilon)c_{\gamma}r^{2\alpha-1}\dfrac{(1+w_{\varepsilon,\pm}^{2})^{\alpha}}{w_{\varepsilon,\pm}^{2\alpha-1}},\>r>0\\ w_{\varepsilon(0),\pm}=0,\mbox{ and }\alpha>\frac{1}{2}.\end{cases}$
which behave as
$w_{\varepsilon,\pm}=\exp\!\Bigl(\frac{(1\pm\varepsilon)c_{\gamma}r^{2\alpha}}{2\alpha}\Bigr)\bigl(1+o(1)\bigr),\mbox{ as }r\to\infty.$
Proposition 4.13.
For $k_{\gamma}=2\alpha-1$ and $\alpha>\frac{1}{2}$, the functions $w_{\varepsilon,+}(r)$ and $w_{\varepsilon,-}$ are supersolution and subsolution, respectively, to Eq. (15) with initial condition $w(0)=0$ for large enough $r$.
Proof.
A computations reveals
$w_{\varepsilon,\pm}^{\prime}-(1+w_{\varepsilon,\pm}^{2})g_{+}\left(\frac{w_{\varepsilon,\pm}}{r(1+w_{\varepsilon,\pm}^{2})},1\right)$
$=\frac{c_{\gamma}r^{k_{\gamma}}}{w_{\varepsilon,\pm}^{k_{\gamma}}}(1+w_{\varepsilon,\pm}^{2})^{\alpha}-(1+w_{\varepsilon,\pm}^{2})g_{+}\left(\frac{w_{\varepsilon,\pm}}{r(1+w_{\varepsilon,\pm}^{2})},1\right)$
$=\pm\dfrac{c_{\gamma}r^{2\alpha-1}}{w_{\varepsilon,\pm}^{2\alpha-1}}(1+w_{\varepsilon,\pm}^{2})^{\alpha}\varepsilon+o\left(r^{2\alpha-1}\dfrac{(1+w_{\varepsilon,\pm}^{2})^{\frac{(2\alpha-1)(\alpha-1)}{2\alpha}}}{w_{\varepsilon,\pm}^{2\alpha-1}}\right).$
Therefore, by the above signs, we obtain that $w_{\varepsilon,+}$ is a supersolution and $w_{\varepsilon,-}$ is a subsolution to equation (15) with initial condition $w_{\varepsilon,\pm}(0)=0$. ∎
Then, by the above proposition and taking $\varepsilon\to 0^{+}$ on $w_{\varepsilon,+}\leq v\leq w_{\varepsilon,+}$, we obtain that $v=e^{c_{\gamma}\frac{r^{2\alpha}}{2\alpha}+o(r^{2})}$, finishing the second part of Thm 4.8.
We end this section by showing an explicit example where this behavior applies. Let $\gamma=\sqrt[n]{S_{n}}$. In this case, $x=g_{+}(y,1)=\frac{1}{y^{n-1}}$ with $n\geq 2$, and thus the translator ODE is given by
$v^{\prime}=(1+v^{2})\left(\frac{r}{v}\right)^{n-1}.$
Therefore, the slope of the bowl-type solution satisfies
$v(r)=\begin{cases}e^{(}\frac{r^{2}}{2}+o(r^{2}),&\mbox{ for }n=2,\\ \left(\frac{n-2}{n}\right)^{\frac{1}{n-2}}r^{\frac{n}{n-2}}+o\left(r^{\frac{n}{n-2}}\right),&\mbox{ for }n\geq 3.\end{cases},\mbox{ as }r\to\infty,$
see figure below.
0112233011223344rrv(r)v(r)n=2n=2n=3n=3n=1500n=1500
Figure 4. Large-$r$ behavior of $v(r)$ for $n=2,3,1500$.